OpenStudy (anonymous):

Which type of discontinuity exists at x = 2 for f(x)=x^2−4/x−2? removable discontinuity jump discontinuity infinite discontinuity none of the above

OpenStudy (michele_laino):

hint: we can factor the numerator as follows: \[\Large {x^2} - 4 = \left( {x - 2} \right)\left( {x + 2} \right)\]

OpenStudy (michele_laino):

factorize*